OpenStudy (anonymous):

y^2*3y^3 is simplified to what?

OpenStudy (anonymous):

3y^5 do you know why?

OpenStudy (anonymous):

i understand the ^5 part. and it would stay 3y because y is basically 1y so when you multiply them, it would become 3y... right..?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you. c:

OpenStudy (unklerhaukus):

\[y^2\times3y^3\]\[=(y\times y)\times3\times(y\times y\times y)\]\[=3\times y\times y\times y\times y\times y\]\[=3y^5\]